OpenStudy (anonymous):

Use natural logarithms to solve the equation. 5e^(2x+11)=30

OpenStudy (solomonzelman):

hit both sides with a log.

OpenStudy (solomonzelman):

with a natural log.

OpenStudy (solomonzelman):

try....please.

OpenStudy (isaiah.feynman):

Lol "hit"

OpenStudy (anonymous):

Alright hold on :P

OpenStudy (solomonzelman):

I remember my trigonometry teacher, (a quite knowledgeable man in math and science) called that "hitting with a log".

OpenStudy (solomonzelman):

take your time, priest.

OpenStudy (isaiah.feynman):

Oh no Zelman, we have to do something first before hitting both sides with a log.

OpenStudy (solomonzelman):

yes, true, lol

OpenStudy (solomonzelman):

Divide both sides by 5.

OpenStudy (isaiah.feynman):

Not what I meant..

OpenStudy (solomonzelman):

stop trying to confuse me, I already suck at math.

OpenStudy (solomonzelman):

Yes, go ahead and divide both side by 5.

OpenStudy (anonymous):

You can actually do it either way, so long as you know the properties of logs. It's just that division is a lower order operation than log, so dividing first is simpler.

OpenStudy (anonymous):

e^(2x+11)=6 :P

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

now take the \(\large\color{black}{ \ln }\) of both sides.

OpenStudy (isaiah.feynman):

I love a neater way, if you are gonna do logarithms first, you can take the log (base 5) of both sides.

OpenStudy (solomonzelman):

or you can apply a rule: \(\large\color{black}{ a=e^{\ln a} }\)

OpenStudy (anonymous):

2x + 11 = ln 6.... Do I just subtract the 11?

OpenStudy (isaiah.feynman):



OpenStudy (solomonzelman):

yes, you are getting it right. 2x+11=ln(6).

OpenStudy (solomonzelman):

and yes again, subtract 11 from both sides.

OpenStudy (anonymous):

2x = ln (-5) Divide 2 from both sides... x = ln(-5)/2